# Observational Features of Thin Accretion Disk Around Rotating Regular Black Hole

### v1

@@ I Introduction
General relativity (GR) as the best theory of gravitational interaction is strongly supported by recent observations of gravitational waves [1, 2, 3] and black hole images [4, 5]. However, it faces the spacetime singularity problem, which is widely considered non-physical and highlights the limitations of general relativity [6, 7]. Alternative black hole without singularity or the so-called regular black hole is a promising way to explore the physics beyond GR. There are many kinds of regular black hole models [8, 9, 10], including the famous Bardeen [11] and Hayward models [12]. Recently, a new regular black hole model attracts the attention of gravitational community, which modifies the black hole mass with an exponential converge factor [13, 14], yielding regular spacetime metrics, including the rotating counterparts [15, 16, 17]. Interestingly, this black hole model has a asymptotic flat spacetime metric in its core, just like the peace in the eye of storm [16]. While some studies have examined their phenomenology, focusing on black hole images, quasinormal modes , energy extraction mechanism and observational testing [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], broader astrophysical tests remain limited and warrant further investigation.
@@
The accretion disk is usually modeled as thin accretion disk. The thin accretion disk model was originally proposed by Novikov-Thorne [38], and Page–Thorne [39], as well as Shakura and Sunyaev [40], assumes that the accretion disk is geometrically thin and optically thick. Under these assumptions, the accretion disk is closely approximate as a blackbody radiation object [38, 39, 40]. This allows for simplified yet highly effective descriptions of disk behavior, making it a widely adopted model for studying accretion processes in various astrophysical scenarios, including black holes. Even recently, there are plenty of works on studying the observational signatures of thin accretion disk in different black hole spacetimes [42, 43, 44, 45, 46, 47, 48, 49, 50], focusing on the effects on the flux, the temperature, luminosity, and so on.
@@
This paper is organized as follows: In Sec.II, we will introduce the rotating regular black hole and discuss some useful quantities of studying the observational features for both time-like and null-like geodesics. In Sec.III, we will study the emission properties of thin accretion disk in this rotating regular spacetime metric, and showing how the regular parameter affects observables of accretion disk. Then, in Sec.IV, we will obtain the realistic black hole image of thin accretion disk based on the ray-tracing method, and comparing the results with Kerr metric and discussing the effects of rotating regular black hole on the lensing bands and the images. In Sec.V, we will make a conclusion and discussion on the results of this work.
@@ II Geodesic properties of rotating regular black hole
$ds^{2}=g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2g_{t\phi}dtd\phi$ (1)
@@
$g_{\phi\phi}$ $=\left(r^{2}+a^{2}+\frac{2Mra^{2}\mathrm{e}^{-k/r}}{\Sigma}\sin^{2}\theta\right)\sin^{2}\theta$ (2)
@@
$\frac{\Sigma}{E}p^{\phi}$ $=\frac{a}{\Delta}\left(r^{2}+a^{2}-a\lambda\right)+\frac{\lambda}{\sin^{2}\theta}-a$ (5)
$\frac{\Sigma}{E}p^{t}$ $=\frac{r^{2}+a^{2}}{\Delta}\left(r^{2}+a^{2}-a\lambda\right)+a\left(\lambda-a\sin^{2}\theta\right)$ (6)
@@
$\mathcal{R}(r)$ $=\left(r^{2}+a^{2}-a\lambda\right)^{2}-\Delta\left[\mu^{2}r^{2}+\eta+(\lambda-a)^{2}\right]$ (7)
$\Theta(\theta)$ $=\eta+a^{2}\cos^{2}\theta-\lambda^{2}\cot^{2}\theta$ (8)
for photons $\mu=0$, while for particles one is free to set $\mu=1$ which indicates the conserved quantities $E$, $\lambda$, and $\eta$ are all normalized with particle mass $\mu$, and we are considering the conserved quantities of time-like geodesics in the per unit mass basis.
@@ II.1 Key quantities of time-like geodesics
We consider that the particles of thin accretion disk are all massive and moving in circular orbits at the equatorial plane, which follow the time-like geodesics. In order to study the observational properties of thin accretion disk, we need first to know several key quantities of the particles in the thin accretion disk. Although we are interested in the particles moving circularly in the equatorial plane where $Q=0$ and $\theta=\pi/2$, we still provide a very general form to the key quantities in the following. The general form is more concise and universal than the specific cases.
@@
$\Omega=\frac{-\partial_{r}g_{t\phi}\pm\sqrt{\left(\partial_{r}g_{t\phi}\right)^{2}-\left(\partial_{r}g_{tt}\right)\left(\partial_{r}g_{\phi\phi}\right)}}{\partial_{r}g_{\phi\phi}}$ (9)
With the expression of $\Omega$, one could obtain the energy and angular momentum per unit particle mass of this test particle as follows,
$E=-\frac{g_{tt}+g_{t\phi}\Omega}{\sqrt{-g_{tt}-2g_{t\phi}\Omega-g_{\phi\phi}\Omega^{2}}}$ (10)
$L=\frac{g_{t\phi}+g_{\phi\phi}\Omega}{\sqrt{-g_{tt}-2g_{t\phi}\Omega-g_{\phi\phi}\Omega^{2}}}$ (11)
@@
$u^{t}=\frac{1}{\sqrt{-g_{tt}-2\Omega g_{t\varphi}-\Omega^{2}g_{\varphi\varphi}}}$ (12)
@@
$\frac{d^{2}\mathcal{R}(r)}{dr^{2}}\bigg{|}_{r_{\text{ISCO}}}\,=0$ (13)
Dose not like the Kerr metric, we do not have an analytical formula for $r_{\text{ISCO}}$, rather we could only compute it numerically.
The $r_{\text{ISCO}}$ for the circular orbits in the equatorial plane with different $k$ and black hole spin $a$ are shown in Fig.1. We can see that as the spin parameter $a$ increases, the radius of $r_{\text{ISCO}}$ decreases monotonically for all $k$ values. This is the same to the behavior predicted by Kerr black hole models, where higher spins lead to more compact stable orbits. For fixed values of $a$, an increase in $k$ leads to an overall decrease in $r_{\text{ISCO}}$. This indicates that the parameter $k$ introduces a deviation from the standard Kerr black hole, effectively decreasing the radius of stable orbits. The $r_{\text{ISCO}}$ is always greater than the event horizon $r_{+}$ as we expected. The larger regular parameter $k$, the smaller the maximum spin. As the spins approach their maximum, $r_{\text{ISCO}}$ coincides with the event horizons of rotating regular black hole.
@@ II.2 Key radii of null-like geodesics
For photons $\mu=0$, it will follow the null-like geodesics. Although it is not related to the composition of thin accretion disk in our assumption, it is crucial to the ray-tracing method which is used to obtain the black hole image. We will discuss several important radii that will be used a lot in the ray-tracing method and thus in the black hole imaging process.
@@
$r^{4}+\mathcal{A}r^{2}+\mathcal{B}r+\mathcal{C}=0$ (14)
@@
$\mathcal{A}$ $=a^{2}-\eta-\lambda^{2}$ (15)
$\mathcal{B}$ $=2Me^{-k/r}\left[\eta+(\lambda-a)^{2}\right]>0$
$\mathcal{C}$ $=-a^{2}\eta$
@@
$\mathcal{R}(r)\bigg{|}_{\tilde{r}}=\frac{d\mathcal{R}(r)}{dr}\bigg{|}_{\tilde{r}}=0$ (16)
For $\tilde{r}>r_{+}$, the conserved quantities $\lambda$, $\eta$ can only take the following unique values
@@
$\tilde{\eta}$ $=\frac{\tilde{r}^{3}}{a^{2}}\left[\frac{4\tilde{\Delta}(M-N\tilde{r})}{(\tilde{r}-M-N\tilde{r})^{2}}-\tilde{r}\right]$ (18)
$N$ $=\frac{k}{\tilde{r}^{2}}e^{-k/\tilde{r}}$ (19)
where $\tilde{\Delta}=\tilde{r}^{2}+a^{2}-2M\tilde{r}e^{-k/\tilde{r}}$. When we set $\tilde{\eta}=0$, we will obtain two radii that greater than the horizon $\tilde{r}_{\pm}>r_{+}$, $\tilde{r}_{-}$ and $\tilde{r}_{+}$ corresponding to the inner and out boundary radius of photon shell, respectively. All the radius of bounded orbits $\tilde{r}$ lie within the range $\tilde{r}\in[\tilde{r}_{-},\,\tilde{r}_{+}]$. Bounded orbits are not stable and will form a critical curve in the observe image plane, see Sec.IV. We have shown the inner edge of photon shell $\tilde{r}_{-}$ for different $k$ and black hole spin $a$ in Fig.1
@@
We can see that as the spin parameter $a$ increases, $\tilde{r}_{-}$ decreases for all values of $k$. For a fixed $a$, the value of $\tilde{r}_{-}$ decreases slightly with increasing $k$. This indicates that the deviation parameter $k$ alters the black hole geometry, pushing the $\tilde{r}_{-}$ inward. The gap between $\tilde{r}_{-}$ and $r_{+}$ becomes narrower as $a$ approaches the maximum, indicating that high-spin black holes exhibit a more compact structure. The presence of $k$ speed up this narrowing, suggesting its role in suppressing the distinction between the two radii. As the spins approach their maximum, $\tilde{r}_{-}$ also coincides with the event horizons of rotating regular black hole.
@@ III Emission properties of thin accretion disks
In thin accretion disk model [38, 39, 40], we have the radiative flux $\mathcal{F}(r)$ which represents the energy emitted per unit area and per unit time at radius $r$ from the accretion disk. It takes form as
@@
where $\dot{m}$ is the mass accretion rate, $g$ is the metric determinant of three dimensional subspace $(t,r,\phi)$
$g=g_{rr}(g_{tt}g_{\phi\phi}-g_{t\phi}^{2})$ (21)
@@
We can see that for a non-rotating black hole ($a=0$), the radiative flux corresponding to different $k$ exhibit minimal differences, indicating weak effects of the regular parameter $k$. However, as the spin parameter increases, the deviations between different $k$ grow significantly, especially the inner region of the thin accretion disk. This behavior shows the combined influence of $a$ and $k$ on the physical properties of the thin accretion disk. The inner radii where radiative flux approach to zero are the corresponding $r_{\text{ISCO}}$. We could see that they are shifted to smaller radius for large black hole spin and regular parameter. The values of radiative flux also increases as the spin and regular parameter increase. In the far region, they tend to be the same regardless the black hole spin $a$ and regular parameter $k$.
@@
Assuming the thin accretion disk is in the state of thermal equilibrium, we could obtain the temperature according to the Stefan-Boltzmann law
$T=\sqrt[4]{\mathcal{F}(r)/\sigma}$ (22)
@@
where $\sigma$ is the Stefan–Boltzmann constant. The distributions of temperature as a function of radial distance with different black hole spin $a$ and regular parameter $k$ are shown in Fig.3. The results of Kerr black hole ($k=0$) are also provided as references. The values in this figure are also multiplied by $10^{5}$.
@@
Figure 3: The temperature of thin accretion disk $T(r)$ as a function of $r/M$ for different black hole spin $a$ and regular parameter $k$.The values in this figure are also multiplied by $10^{5}$.
@@
From Fig.3, we know that as the spin parameter increases, the temperature distributions show larger deviations, particularly at the inner region of the accretion disk. The larger black hole spin $a$ and regular parameter $k$, the higher temperature of the accretion disk which means greater radiation efficiency. The inner radii where temperature approach to zero are also the corresponding $r_{\text{ISCO}}$. At larger radii, the temperature curves converge, reflecting the decreasing influence of both black hole spin $a$ and regular parameter $k$ in the far region.
@@
Another interesting quantity we wish to study is the so-called differential luminosity which could be measured by observations. It takes the form as
$\frac{d\mathcal{L}_{\infty}}{d\ln r}=4\pi r\sqrt{-g}E\mathcal{F}(r)$ (23)
@@
where $\mathcal{L}_{\infty}$ represents the energy per unit of time reaching an observer at infinity. We have shown the differential luminosity as a function of radial distance with different black hole spin $a$ and regular parameter $k$ in Fig.4. The values in this figure are also multiplied by $10^{5}$ for convenience.
@@
Figure 4: Differential luminosity of thin accretion disk $d\mathcal{L}_{\infty}/d\ln r$ as a function of $r/M$ for different black hole spin $a$ and regular parameter $k$. The results of Kerr black hole ($k=0$) are also provided as references. The values in this figure are also multiplied by $10^{5}$.
@@
Again, for slowly rotating black holes ($a=0$), the differential luminosity for different $k$ values are nearly identical, suggesting a weak dependence on the regular parameter at low black hole spin. As the spin increases, the differential luminosity changes more significantly with the regular parameter $k$, particularly in the region near the black hole. The larger black hole spin $a$ and regular parameter $k$, the larger overall differential luminosity. The inner radii where differential luminosity approach to zero are also the corresponding $r_{\text{ISCO}}$. At larger radii, all curves converge, indicating that the effects of $a$ and $k$ diminish in the far region.

### v2

@@ I Introduction
General relativity (GR) as the best theory of gravitational interaction is strongly supported by recent observations of gravitational waves [1, 2, 3] and black hole images [4, 5]. However, it faces the spacetime singularity problem, which is widely considered non-physical and highlights the limitations of general relativity [6, 7]. Alternative black hole without singularity or the so-called regular black hole is a promising way to explore the physics beyond GR. There are many kinds of regular black hole models [8, 9, 10], including the famous Bardeen [11] and Hayward models [12]. Recently, a new regular black hole model attracts the attention of gravitational community, which modifies the black hole mass with an exponential converge factor [13, 14] (see also a geodesically complete version [15]), yielding regular spacetime metrics. The rotating counterparts of this regular black hole are also obtained in [16, 17, 18]. Interestingly, this black hole model has a asymptotic flat spacetime metric in its core, just like the peace in the eye of storm [17]. While some studies have examined their phenomenology, focusing on black hole images, quasinormal modes, energy extraction mechanism, primordial black holes, and observational tests [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], broader astrophysical tests remain limited and warrant further investigation.
@@
The accretion disk is usually modeled as thin accretion disk, which was originally proposed by Novikov-Thorne [40], and Page–Thorne [41], as well as Shakura and Sunyaev [42], assuming that the accretion disk is geometrically thin and optically thick. Under these assumptions, the accretion disk radiates approximately as a blackbody at each radius [40, 41, 42]. This allows for simplified yet highly effective descriptions of disk behavior, making it a widely adopted model for studying accretion processes in various astrophysical scenarios, including black holes. Even recently, there are plenty of works on studying the observational signatures of thin accretion disk in different black hole spacetimes [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], focusing on the effects on the flux, the temperature, luminosity, images, and so on.
@@
This paper is organized as follows: In Sec. II, we will introduce the rotating regular black hole and discuss some useful quantities or radii of studying the observational features for both time-like and null-like geodesics. In Sec. III, we will study the emission properties of thin accretion disk in this rotating regular spacetime, and showing how the regular parameter affects observables of accretion disk. Then, in Sec. IV, we will obtain the realistic bolometric image of thin accretion disk based on the ray-tracing method, and comparing the results with Kerr black hole, discussing the effects of rotating regular black hole on the lensing bands and the images. In Sec. V, we will make a conclusion and discussion on the results of this work.
@@ II Geodesic properties of rotating regular black hole
$ds^{2}=g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2g_{t\phi}dtd\phi,$ (1)
@@
$g_{\phi\phi}$ $=\left(r^{2}+a^{2}+\frac{2Mra^{2}\mathrm{e}^{-k/r}}{\Sigma}\sin^{2}\theta\right)\sin^{2}\theta,$ (2)
@@
$\frac{\Sigma}{E}p^{\phi}$ $=\frac{a}{\Delta}\left(r^{2}+a^{2}-a\lambda\right)+\frac{\lambda}{\sin^{2}\theta}-a,$ (5)
$\frac{\Sigma}{E}p^{t}$ $=\frac{r^{2}+a^{2}}{\Delta}\left(r^{2}+a^{2}-a\lambda\right)+a\left(\lambda-a\sin^{2}\theta\right),$ (6)
@@
$\mathcal{R}(r)$ $=\left(r^{2}+a^{2}-a\lambda\right)^{2}-\Delta\left[\mu^{2}r^{2}+\eta+(\lambda-a)^{2}\right],$ (7)
$\Theta(\theta)$ $=\eta+a^{2}\cos^{2}\theta-\lambda^{2}\cot^{2}\theta.$ (8)
For photons $\mu=0$. While for massive particles, one is free to set $\mu=1$, which indicates the conserved quantities $E$, $\lambda$, and $\eta$ are all normalized with particle mass $\mu$. We are considering the conserved quantities of time-like geodesics in the per unit mass basis.
@@ II.1 Key quantities of time-like geodesics
We assume that the particles composing the thin accretion disk are all massive and moving in circular orbits at the equatorial plane, which follow the time-like geodesics. In order to study the observational properties of thin accretion disk, we need first to know several key quantities of these particles. Although we are interested in the particles moving circularly in the equatorial plane where $Q=0$ and $\theta=\pi/2$, we still provide a very general form to the key quantities in the following. The general form is more concise and universal than the specific cases.
@@
$\Omega=\frac{-\partial_{r}g_{t\phi}\pm\sqrt{\left(\partial_{r}g_{t\phi}\right)^{2}-\left(\partial_{r}g_{tt}\right)\left(\partial_{r}g_{\phi\phi}\right)}}{\partial_{r}g_{\phi\phi}}.$ (9)
With the expression of $\Omega$, one could obtain the energy and angular momentum per unit particle mass of a test particle as follows,
$E=-\frac{g_{tt}+g_{t\phi}\Omega}{\sqrt{-g_{tt}-2g_{t\phi}\Omega-g_{\phi\phi}\Omega^{2}}},$ (10)
$L=\frac{g_{t\phi}+g_{\phi\phi}\Omega}{\sqrt{-g_{tt}-2g_{t\phi}\Omega-g_{\phi\phi}\Omega^{2}}},$ (11)
@@
$u^{t}=\frac{1}{\sqrt{-g_{tt}-2\Omega g_{t\varphi}-\Omega^{2}g_{\varphi\varphi}}}.$ (12)
@@
$\frac{d^{2}\mathcal{R}(r)}{dr^{2}}\bigg{|}_{r_{\text{ISCO}}}\,=0.$ (13)
Dose not like the Kerr metric, we do not have an analytical formula for $r_{\text{ISCO}}$, rather we could only compute it numerically. We adopted the numerical method from [58].
The $r_{\text{ISCO}}$ for the circular orbits in the equatorial plane with different $k$ and black hole spin $a$ are shown in Fig. 1. We can see that as the spin parameter $a$ increases, the radius of $r_{\text{ISCO}}$ decreases monotonically for all $k$ values. This is the same to the behavior predicted by Kerr black hole models, where higher spins lead to more compact stable orbits. For fixed values of $a$, an increase in $k$ leads to an overall decrease in $r_{\text{ISCO}}$. This indicates that the parameter $k$ introduces a deviation from the standard Kerr black hole, effectively decreasing the radius of stable orbits. The $r_{\text{ISCO}}$ is always greater than the event horizon $r_{+}$ as we expected. The $r_{\text{ISCO}}$ and event horizon will become coincident as the spin reaches a maximum value. When beyond the maximum spin, the rotating regular black hole will have no horizons. The larger regular parameter $k$, the smaller the maximum spin.
@@ II.2 Key radii of null-like geodesics
For photons $\mu=0$, it will follow the null-like geodesics. Although it is not related to the composition of thin accretion disk, it is crucial to the ray-tracing method, which is used to obtain the black hole image. We will discuss several important radii that will be used a lot in the ray-tracing method and thus in the black hole imaging process.
@@
$r^{4}+\mathcal{A}r^{2}+\mathcal{B}r+\mathcal{C}=0,$ (14)
@@
$\mathcal{A}$ $=a^{2}-\eta-\lambda^{2},$ (15)
$\mathcal{B}$ $=2Me^{-k/r}\left[\eta+(\lambda-a)^{2}\right]>0,$
$\mathcal{C}$ $=-a^{2}\eta.$
@@
$\mathcal{R}(r)\bigg{|}_{\tilde{r}}=\frac{d\mathcal{R}(r)}{dr}\bigg{|}_{\tilde{r}}=0.$ (16)
For $\tilde{r}>r_{+}$, to fulfill the above equation, the conserved quantities $\lambda$, $\eta$ can only take the following unique values
@@
$\tilde{\eta}$ $=\frac{\tilde{r}^{3}}{a^{2}}\left[\frac{4\tilde{\Delta}(M-N\tilde{r})}{(\tilde{r}-M-N\tilde{r})^{2}}-\tilde{r}\right],$ (18)
$N$ $=\frac{k}{\tilde{r}^{2}}e^{-k/\tilde{r}},$ (19)
where $\tilde{\Delta}=\tilde{r}^{2}+a^{2}-2M\tilde{r}e^{-k/\tilde{r}}$. In the equatorial plane, we have $\tilde{\eta}=0$, then this condition will provide us two radii that greater than the horizon $\tilde{r}_{\pm}>r_{+}$, $\tilde{r}_{-}$ and $\tilde{r}_{+}$ corresponding to the inner and out boundary radius of photon shell, respectively. All the radius of bounded orbits $\tilde{r}$ lie within the range $\tilde{r}\in[\tilde{r}_{-},\,\tilde{r}_{+}]$. Bounded orbits are not stable and will form a critical curve in the observer’s image plane, see Sec. IV. We have shown the inner edge of photon shell $\tilde{r}_{-}$ for different $k$ and black hole spin $a$ in Fig. 1
@@
We can see that as the spin parameter $a$ increases, $\tilde{r}_{-}$ decreases for all values of $k$. For a fixed $a$, the value of $\tilde{r}_{-}$ decreases slightly with increasing $k$. This indicates that the deviation parameter $k$ alters the black hole geometry, pushing the $\tilde{r}_{-}$ inward. The gap between $\tilde{r}_{-}$ and $r_{+}$ becomes narrower as $a$ approaches the maximum, indicating that high-spin black holes exhibit a more compact structure. The presence of $k$ speed up this narrowing, suggesting its role in suppressing the distinction between the two radii. As the spins approach their maximum, $\tilde{r}_{-}$ also coincides with the event horizon, the same to $r_{\text{ISCO}}$, of the rotating regular black hole.
@@ III Emission properties of thin accretion disks
In thin accretion disk model [40, 41, 42], we have the radiative flux $\mathcal{F}(r)$ which represents the energy emitted per unit area and per unit time at radius $r$ from the thin accretion disk. It takes form as
@@
where $\dot{m}$ is the mass accretion rate, and here we consider $\dot{m}$ is much smaller than the Eddington accretion rate, where the thin accretion disk model is valid. $g$ is the metric determinant of three dimensional subspace $(t,r,\phi)$
$g=g_{rr}(g_{tt}g_{\phi\phi}-g_{t\phi}^{2}).$ (21)
@@
We can see that for a non-rotating black hole ($a=0$), the radiative flux corresponding to different $k$ exhibit minimal differences, indicating weak effects of the regular parameter $k$. However, as the spin parameter increases, the deviations between different $k$ grow significantly, especially the inner region of the thin accretion disk. The larger $k$, the lager radiative flux, which means greater radiation efficiency. The figure also shows the combined influence of $a$ and $k$ on the physical properties of the thin accretion disk. The inner radii where radiative flux approach to zero are the corresponding $r_{\text{ISCO}}$. We could see that they are shifted to smaller radius for large black hole spin and regular parameter. The values of radiative flux also increases as the spin and regular parameter increase. In the far region, they tend to be the same regardless the black hole spin $a$ and regular parameter $k$.
@@
Assuming at each radius $r$, the thin accretion disk is in the state of thermal equilibrium, we could then obtain the temperature according to the Stefan-Boltzmann law
$T=\sqrt[4]{\mathcal{F}(r)/\sigma},$ (22)
@@
where $\sigma$ is the Stefan–Boltzmann constant. The distributions of temperature as a function of radial distance with different black hole spin $a$ and regular parameter $k$ are shown in Fig. 3. The results of Kerr black hole ($k=0$) are also provided as references. The Stefan–Boltzmann constant $\sigma$ is set to be 1, and the values in this figure are all multiplied by $10^{2}$ for convenience.
@@
Figure 3: The temperature of thin accretion disk $T(r)$ as a function of $r/M$ for different black hole spin $a$ and regular parameter $k$. We set the Stefan–Boltzmann constant $\sigma=1$, and the values in this figure are multiplied by $10^{2}$ for convenience
@@
From Fig. 3, we know that as the spin parameter increases, the temperature distributions show larger deviations, particularly at the inner region of the accretion disk. The larger black hole spin $a$ and regular parameter $k$, the higher temperature of the thin accretion disk. The inner radii where temperature approach to zero are also the corresponding $r_{\text{ISCO}}$. At larger radii, the temperature curves converge, reflecting the decreasing influence of both black hole spin $a$ and regular parameter $k$ in the far region.
@@
Another interesting quantity we wish to study is the so-called differential luminosity of the thin accretion disk, which could be measured by observations. It takes the form as
$\frac{d\mathcal{L}_{\infty}}{d\ln r}=4\pi r\sqrt{-g}E\mathcal{F}(r),$ (23)
@@
where $\mathcal{L}_{\infty}$ represents the energy per unit of time reaching an observer at infinity. We have shown the differential luminosity as a function of radial distance with different black hole spin $a$ and regular parameter $k$ in Fig. 4. The values in this figure are multiplied by $10^{2}$ for convenience.
@@
Figure 4: Differential luminosity of thin accretion disk $d\mathcal{L}_{\infty}/d\ln r$ as a function of $r/M$ for different black hole spin $a$ and regular parameter $k$. The results of Kerr black hole ($k=0$) are also provided as references. The values in this figure are multiplied by $10^{2}$.
@@
For slowly rotating regular black holes, the differential luminosity for different $k$ values are nearly identical, suggesting a weak dependence on the regular parameter at low black hole spin. As the spin increases, the differential luminosity changes more significantly with the regular parameter $k$, particularly in the region near the black hole. The larger black hole spin $a$ and regular parameter $k$, the larger overall differential luminosity. The inner radii where differential luminosity approach to zero are also the corresponding $r_{\text{ISCO}}$. At larger radii, all curves converge, indicating that the effects of $a$ and $k$ diminish in the far region.